OpenStudy (anonymous):

Solve this equation by completing the square. 3x2 - 4x = -2

OpenStudy (anonymous):

3x^2-4x=-2 3x^2=4x+2=0 (x )(x )=0

OpenStudy (anonymous):

(3x )( x )=0

OpenStudy (anonymous):

i know i got to (3x ) (x )=0 but i tried all the numbers i could think of in the spaces

OpenStudy (anonymous):

use this formula [(-b)+- sqr(b^2-4ac)]/2a

OpenStudy (anonymous):

b=-4 a=3 c=2

OpenStudy (anonymous):

i dont understand how to input my equation into the one you gave me. which parts are A,B,and C

OpenStudy (anonymous):

[4 +- sqr(16-4*3*2)]/6 [4+-sqr(16-24)/]6 [4+-aqr(-8)]/6 2/3+-2/6*sqr(-2)

OpenStudy (anonymous):

A: x^2 coefficient B: x coefficient C: the number

OpenStudy (anonymous):

okay thank you a ton

OpenStudy (anonymous):

the sqr of negative number irrational, then the two roots are a conjugate pair.

OpenStudy (anonymous):

YW